OpenStudy (anonymous):

HELP! f(x)=t-2/3t+7 for 1/f(x)

OpenStudy (anonymous):

@Matthew071

OpenStudy (john_es):

Do yo mean, \[f^{-1}(x)\] or \[\frac{1}{f(x)}\] ?

OpenStudy (anonymous):

the 2nd one

OpenStudy (john_es):

Well, then you only have to put it this way, \[\frac{1}{t-\frac{2}{3t}+7}\]I'm not sure if you write 2/(3t) or 2/3*t.

OpenStudy (john_es):

And I think this is a strange problem, asking to find that.

hartnn (hartnn):

i think its \(\large \dfrac{t-2}{3t+7}\) can we have the entire question ??

OpenStudy (anonymous):

i am confused :@

OpenStudy (anonymous):

what i have posted is the entire question.

OpenStudy (john_es):

If @hartnn is correct, then, they are asking you for the inverse function, so it is, \[f^{-1}(x)\]

hartnn (hartnn):

"for 1/f(x)" <<<<<<is not a question.....

OpenStudy (john_es):

True.

OpenStudy (anonymous):

they have given me this function f(x)= t-2/3t+7 and said to find 1/f(x)

OpenStudy (anonymous):

i mean f(t)= t-2/3t+7

OpenStudy (john_es):

But must be as @hartnn wrote it, f(t)= (t-2)/(3t+7)

OpenStudy (anonymous):

yeah

hartnn (hartnn):

to get 1/f(t) you just flip the numerator and denom. \(\large \dfrac{f(t)}{1}=\dfrac{t-2}{3t+7} \\ \large \implies \dfrac{1}{f(t)}=\dfrac{3t+7}{t-2}\) thats it!

OpenStudy (anonymous):

that is exactly what i did, but i was confused whether it was right or wrong :/

hartnn (hartnn):

it is correct.

OpenStudy (anonymous):

thanks :)

hartnn (hartnn):

welcome ^_^